OpenStudy (anonymous):

Write each equation in slope-interept form then identify the slope and y intercept for each line. 2x - y =12

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

This is the only information that I was given for the question I don't know how to put in slope intercept form struggling with this

OpenStudy (noseboy908):

slope intercept from is "y=mx+b", where m= slope, and b = the y-intercept. You have to get y on it's own to have this equation in slope-intercept form.

OpenStudy (anonymous):

so it would be y = 12 - 2x

OpenStudy (anonymous):

Does anyone know if I got the equation correct y = 12-2x

OpenStudy (noseboy908):

show me the steps you used to get to that answer. by the way, the slope always comes before the intercept.

OpenStudy (anonymous):

I subtracted 2x from both sides to y by itself

OpenStudy (noseboy908):

ok, once you do that is should be -y=2x-+12

OpenStudy (noseboy908):

and you cannot have a -y. it has to be positive.

OpenStudy (anonymous):

you lost me

OpenStudy (noseboy908):

ok. your staring equation is "2x-y=12" Then you subtracted 2x from both sides... : "-y=-2x+12" But, you cannot have a -y, it has to be positive. so, divide everything by -1 so that y can be positive. That is the slope-intercept equation.

OpenStudy (anonymous):

ok got it thank you

OpenStudy (noseboy908):

so. what do you get as your slope and y-intercept?

OpenStudy (anonymous):

y = 2x/2 +12/2 y = x + 6 x + 6 = 2(x + 6) + 12 x +6 = 2x +24 6 = 2x/x +24 6 -24 =2x -18/2 = 2x/2 x = 9 y = 2(9) + 12 y = 18 +12 y = 30

OpenStudy (noseboy908):

uh, no. you really over thought it. The slope is whatever is before x, and the y-intercept is the number after the slope. Example: with the equation "y=3x+5", the slope would be 3 (or 3/1, same thing), and the intercept would be 5.